Circular No. 8143

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

1999 RZ_{253}

K. S. Noll and D. C. Stephens, Space Telescope Science Institute, report the detection, with the Hubble Space Telescope (+ camera 2 of the Near-Infrared Camera Multi-Object Spectrometer) on Apr. 23.101–23.181 UT, of a binary companion to the transneptunian object 1999 RZ_{253} (cf. MPEC 2000-O12, 2000-O18; MPO 13608), during an ongoing observing program that includes team members D. Cruikshank, W. Grundy, W. Romanishin, and S. Tegler. The images were obtained through the F110W and F160W filters, with one exposure per filter at each of two dithered positions on the detector, giving two exposures in each filter. Two components are clearly resolved in each image, separated by an angular distance of 0′′.21 ± 0′′.02 at a position angle of 116^{o} ± 4^{o}. The spacecraft tracked the motion of 1999 RZ_{253} as it moved > 4′′.5; the relative position of the two components remained the same during that time. As 1999 RZ_{253} was then 41.55 AU from the earth, the projected separation of the objects in the plane of the sky is 6300 ± 600 km. Earlier observations of this object on 2001 Nov. 9 and 12 with the Space Telescope Imaging Spectrograph show ambiguous evidence for the binary. The images taken on Nov. 9 show inconclusive evidence of a possible unresolved binary at a position angle of 15^{o} ± 15^{o}, while those from Nov. 12 do not show evidence for a second object (the separation of the components on both dates being < 0′′.1).

SUPERNOVAE 2003fb AND 2003fc M. Papenkova and W. Li report the LOTOSS/KAIT discovery of two

additional apparent supernovae (cf. IAUC 8141):

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003fb June 5.5 20^{h}11^{m}50.^{s}33 + 5^{o}45′37′′.6 18.6 14′′.5 E, 12′′.6 S
2003fc June 5.5 20 08 22.08 −17 36 30.5 18.4 2′′.2 W, 7′′.4 N

Additional magnitudes: SN 2003fb in UGC 11522, 2002 Nov. 13.3 UT, [19.0; 2003 June 6.4, 18.6. SN 2003fc, 2002 Oct. 20.3, [19.0; 2003 June 6.4, 18.3. B. Schmidt and M. Salvo, Research School of Astronomy and Astrophysics, Australian National University (ANU); and A. Ford, Monash University, report that a low-signal-to-noise-ratio spectrum (range 350–900 nm) of SN 2003fb, obtained on June 6.7 UT with the ANU 2.3-m telescope (+ DBS spectrograph), shows it to be a type-II supernova, based on the presence of a developed Hα P-Cyg line.

2003 June 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green